OpenStudy (anonymous):

A retangle table is four times as long as it is wide, If it is 196ft^2 what is the lenght and the wideth please

OpenStudy (anonymous):

4w*w=196

OpenStudy (anonymous):

l * w = a

OpenStudy (anonymous):

196/4 = w^2

OpenStudy (anonymous):

sqrt 196/4 = w

OpenStudy (anonymous):

do the arithmetic

OpenStudy (anonymous):

49

OpenStudy (anonymous):

7

OpenStudy (anonymous):

w = 7 and l equals 4 times that which is 28

OpenStudy (anonymous):

by the rule of the area, L * W multiply 28 times 7 and tell me whatyo uget

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

196

OpenStudy (anonymous):

yw